# Re-accretion of Giant Impact Ejecta Can Drive Significant Atmospheric Erosion on Terrestrial Planets

### v1

@@ 4.2 Effect of/on Neighbouring Planets
To emulate the influence of these neighbouring planets in our analysis, we consider a simplified model where the debris mass re-accreted by Earth is reduced by a factor of 3 and evaluate its consequence on atmospheric erosion  [^5] Note that the factor of 3 reduction in accreted particles found by Jackson & Wyatt (2012) may not directly correspond to a proportional decrease in the total mass accreted because the perturbations from neighbouring planets affect the debris at the late stages of evolution, at which point a substantial portion of the debris mass has already been depleted due to collisional grinding. Therefore, this factor of three reduction in mass accretion should be treated as a conservative lower bound. . The resulting atmospheric erosion is shown in Figure 6 as a function of initial atmospheric mass considering ${f_{\rm{esc}}}=0.01$, ${D_{\rm{max}}}=100\ {\rm{km}}$, and $\mu=29$. As expected, a reduced accretion leads to reduced atmospheric losses, particularly for planets with substantial initial atmospheres. In this scenario, Earth experiences $\gtrsim 20\%$ atmospheric erosion for ${m_{\rm{atm0}}}\lesssim 10\ {m_{\rm{atm\earth}}}$, and complete atmospheric erosion occurs for ${m_{\rm{atm0}}}\lesssim{m_{\rm{atm\earth}}}$ on a relatively longer timescale ($\sim 30$ Myr). Complete erosion of more massive atmospheres is likely for larger $f_{\rm{esc}}$} or $D_{\rm{max}}$}. Since Venus is also expected to accrete comparable amounts of debris as Earth (Jackson & Wyatt, 2012), we use our model to estimate the effects of this accretion on the Venusian atmosphere. Given that Venus has a significantly more massive atmosphere (${m_{\rm{atm,Venus}}}/{m_{\rm{atm\earth}}}\simeq 94$), it loses only $2.6\%$ of its atmosphere due to debris accretion. However, the accretion of such large amounts of debris would result in the deposition of a several-km-thick layer of material with Earth-Moon origin. On the other hand, although Mars is estimated to accrete only $1.45\times 10^{-5}\ {M_{\earth}}$, approximately 1/ 671/\ 67th of debris mass accreted by Earth (Jackson & Wyatt, 2012), this is still three orders of magnitude larger than its present-day atmosphere. As a result, even this limited debris accretion would have been sufficient to entirely erode Mars’ present-day atmosphere and potentially remove up to three times its current atmospheric mass. This implies that, if it released substantial non-vaporised debris, Mars must have obtained its atmosphere since the Moon-forming impact or started with a much more massive atmosphere. However, the observed Xenon depletion and isotope fractionation in the modern Martian atmosphere support a impact driven atmospheric evolution scenario, favouring the former possibility (Shorttle et al., 2024).
@@ 4.4 Geochemical Implications
Given the solar system terrestrial planets likely formed via series of giant impacts (e.g., Clement et al., 2024; Raymond & Morbidelli, 2022), repeated episodes of re-accretion could modulate the redox state of Earth over time. Fully accounting for the origin, size distribution and composition, consequences for Earth of its re-accreted small debris are beyond the scope of the paper. However, this cumulative effect may have played a critical role in the establishment of habitable conditions, particularly in the presence of early oceans that could act as oxidizing agents.
@@ Acknowledgements
T.G. gratefully acknowledges support from the Leverhulme Centre for Life in the Universe at the University of Cambridge.

### v2

@@ 4.2 Effect of/on Neighbouring Planets
To emulate the influence of these neighbouring planets in our analysis, we consider a simplified model where the debris mass re-accreted by Earth is reduced by a factor of 3 and evaluate its consequence on atmospheric erosion  [^5] Note that the factor of 3 reduction in accreted particles found by Jackson & Wyatt (2012) may not directly correspond to a proportional decrease in the total mass accreted because the perturbations from neighbouring planets affect the debris at the late stages of evolution, at which point a substantial portion of the debris mass has already been depleted due to collisional grinding. Therefore, this factor of three reduction in mass accretion should be treated as a conservative lower bound. . The resulting atmospheric erosion is shown in Figure 6 as a function of initial atmospheric mass considering ${f_{\rm{esc}}}=0.01$, ${D_{\rm{max}}}=100\ {\rm{km}}$, and $\mu=29$. As expected, a reduced accretion leads to reduced atmospheric losses, particularly for planets with substantial initial atmospheres. In this scenario, Earth experiences $\gtrsim 20\%$ atmospheric erosion for ${m_{\rm{atm0}}}\lesssim 10\ {m_{\rm{atm\earth}}}$, and complete atmospheric erosion occurs for ${m_{\rm{atm0}}}\lesssim{m_{\rm{atm\earth}}}$ on a relatively longer timescale ($\sim 30$ Myr). Complete erosion of more massive atmospheres is likely for larger $f_{\rm{esc}}$} or $D_{\rm{max}}$}. Since Venus is also expected to accrete comparable amounts of debris as Earth (Jackson & Wyatt, 2012), we use our model to estimate the effects of this accretion on the Venusian atmosphere. Given that Venus has a significantly more massive atmosphere (${m_{\rm{atm,Venus}}}/{m_{\rm{atm\earth}}}\simeq 94$), it loses only $2.6\%$ of its atmosphere due to debris accretion. However, the accretion of such large amounts of debris would result in the deposition of a several-km-thick layer of material with Earth-Moon origin. On the other hand, although Mars is estimated to accrete only $1.45\times 10^{-5}\ {M_{\earth}}$, approximately 1/ 671/\ 67th of debris mass accreted by Earth (Jackson & Wyatt, 2012), this is still three orders of magnitude larger than its present-day atmosphere. As a result, even this limited debris accretion would have been sufficient to entirely erode Mars’ present-day atmosphere and potentially remove up to three times its current atmospheric mass. This implies that, if it released substantial non-vaporised debris, Mars must have obtained its atmosphere since the Moon-forming impact or started with a much more massive atmosphere (e.g., Joiret et al., 2025). However, the observed Xenon depletion and isotope fractionation in the modern Martian atmosphere support a impact driven atmospheric evolution scenario, favouring the former possibility (Shorttle et al., 2024).
@@ 4.4 Geochemical Implications
Given the solar system terrestrial planets likely formed via series of giant impacts (e.g., Quintana et al., 2016; Raymond & Morbidelli, 2022; Clement et al., 2024), repeated episodes of re-accretion could modulate the redox state of Earth over time. Fully accounting for the origin, size distribution and composition, consequences for Earth of its re-accreted small debris are beyond the scope of the paper. However, this cumulative effect may have played a critical role in the establishment of habitable conditions, particularly in the presence of early oceans that could act as oxidizing agents.
@@ Acknowledgements
We thank the anonymous referee for insightful comments and constructive suggestions. T.G. gratefully acknowledges support from the Leverhulme Centre for Life in the Universe at the University of Cambridge through the Joint Collaborations Research Project Grant G112026, LGAG/393.